# Vendor-Aware Industrial Agents: RAG-Enhanced LLMs for Secure On-Premise PLC Code Generation

### v1

@@ V Discussion and Industrial Implications
Industrial practitioners prioritize data protection, security, and quality over raw performance [Brehme2025]. This suggests modest performance trade-offs between cloud and local models are acceptable for eliminating external dependencies in security-critical environments. Preliminary observations show local models avoid API overhead, enabling faster iteration cycles critical for interactive development.
While Mitsubishi Electric-focused, the architecture generalizes to other ST dialects via vendor-specific documentation, prompt templates encoding platform constraints, and vendor compiler integration. This means that the solution could be scaled to more CPU series apart from the used iQR series.
Limitations include Mitsubishi Electric-only scope (though architectural components transfer), task coverage and ”specialist” local model behavior with occasional misinterpretation. Human evaluation from a single organizational context may limit generalizability; longitudinal deployment studies are needed to assess adoption barriers and real-world productivity impacts. The local model limits the available context and thus shortens the system prompt and synthetic training data may have downsides not yet visible.

### v2

@@ V Discussion and Industrial Implications
Industrial practitioners prioritize data protection, security, and quality over raw performance [3]. This suggests modest performance trade-offs between cloud and local models are acceptable for eliminating external dependencies in security-critical environments. Preliminary observations show local models avoid API overhead, enabling faster iteration cycles critical for interactive development.
While this work focuses exclusively on Mitsubishi Electric’s iQ-R series, the architectural components are conceptually transferable to other ST implementations. However, we have not empirically validated cross-vendor generalization. Each platform presents distinct function block semantics and data that would require dedicated adaptation efforts. We therefore limit our claims to the Mitsubishi Electric ecosystem.
Limitations include said Mitsubishi Electric-only scope, task coverage and ”specialist” local model behavior with occasional misinterpretation. Human evaluation from a single organizational context may limit generalizability; longitudinal deployment studies are needed to assess adoption barriers and real-world productivity impacts. The local model limits the available context and thus shortens the system prompt and synthetic training data may have downsides not yet visible.
We acknowledge that compilation success is a necessary but not sufficient criterion for code quality. Functional correctness would require domain-expert evaluation or execution on physical hardware, which we identify as a limitation of this study even though we let expert users validate compilation results qualitatively. Future work should incorporate simulation-based testing or hardware-in-the-loop validation to assess runtime behavior.
Automatically generated PLC code for industrial control systems poses inherent risks that must be addressed in deployment contexts. Faulty control logic could lead to equipment damage, production downtime, or safety hazards for personnel. We emphasize that our system is designed as an assistive tool, not a replacement for human verification. All generated code must undergo a mandatory review by qualified automation engineers before deployment or simulation testing in vendor-provided environments. Moreover, it must undergo compliance verification against site-specific safety standards (e.g., cyber security).